Fermi National Laboratory
 

Operation Summary for October 26, 2009 - October 28, 2009

 

Monday Oct. 26

The day shift began with Operations monitoring the Tevatron (TeV) store 7292 (102.73E30) and the Recycler stash (326.1E10), with the Antiproton Source stacking (28.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7292 at 11:16 AM. FESS personnel cleaned the TeV sector F0 pond water strainer. Cryo system personnel worked on the CHL helium compressors and brought them back on line. Operations put the TeV into HEP shot setup and established store 7293 at 1:49 PM with an average initial luminosity (AIL) of 282.23E30. At 2 PM, Operations halted stacking so Pbar personnel could conduct stacktail studies. Controls personnel worked on the vacuum crate at MI-30. NuMI personnel accessed MI-31. Pbar resumed stacking at 3:14 PM and NuMI resumed taking beam.

During the evening shift, MI personnel investigated beam position problems. A Booster RF station (BRF18) tripped off three times during the shift; it reset. Booster personnel tuned.

Tuesday Oct. 27

The midnight shift began with Operations monitoring store 7293 (103.0 E30) and the stash (235.2E10), with Pbar stacking (21.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported short downtimes due to the following device trips: NuMI BPMs, BRF18, LRF3, I:Q703, and a Linac Klystron RF station (KRF3). Operations terminated store 7293 at 7:27 AM and put the TeV into shot setup.

During the day shift, Operations established store 7294 at 9:17 AM with a luminosity of 272.14E30 per D0; the CDF luminosity counter had a problem. Operations halted stacking for an hour so Pbar personnel could investigate a stacktail problem. The MI Power Supply interlock Loop tripped off at 1:05 PM – operators and MI personnel investigated; it reset.

During the evening shift, the Tevatron Electron Lens (TEL1) tripped off several times on a vacuum fault; it reset. Operators tuned the Booster.

Wednesday Oct. 28

The midnight shift began with Operations monitoring store 7294 (66.59E30) and the stash (234.7E10), with Pbar stacking (17.7E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, CDF experimenters calibrated their luminosity counter. The P1 south transfer line tripped off due to a power supply (I:HT702) wandering high; it reset.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow MI, Pbar, Recycler, and TeV studies.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/28/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory